A spark-ignition engine operates on an Otto cycle with a compression ratio of 9 and temperature limits of 30 degrees C and 1000 degrees C. If the power output is 500 kw, calculate the thermal efficiency and the mass flux of air. (Answers: 58.5%, 2.19 kg/s -- Solution steps needed!!)
TEXTBOOK: Schaum's Thermodynamics for Engineers (Second Edition)
1
Expert's answer
2012-11-12T11:07:49-0500
η = 1- (1/r^(k-1));
r is the compression ratio
k (CP/ Cv) [for air k=1.4]
η = 1- (1/9^(1.4-1))
η = 1- (1/9^0.4);
η = 1- (1/2.41)
η = 1-0.415
η =0.585 or 58.5 %